40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3264     Entry time: Thu Jul 22 03:08:27 2010
Author: Koji 
Type: Update 
Category: IOO 
Subject: Vent and MC lock 

Summary

- The vacuum chambers have been vented.

- The north heavy door of the BS chamber has been opened by Genie (not by the crane).
It was replaced by the light door. The door is currently closed.

- The MC has been locked with 20mW incident and aligned. MC REFL was left unchanged but lock was able to be achieved.

- The optics before MCT CCD and MCT QPD have been adjusted for the low power operation.


Details

- The first HWP for the variable optical attanuator (HWP/PBS/HWP pair) was set to be 86deg from the maximum transmission at 126deg.
The incident power of 19mW has been measured.

- The PSL mechanical shutter has been manually opened. Two other beam blocks has been removed.

- I found the MC was totaly misaligned with no resonance.- I tried to align it based on the previous OSEM values but in vain.

How to align the MC mirrors from the scratch

- MC1 has been aligned so as to maximize REFL PD and DC signal of WFS QPDs.

- MC3 has been aligned by looking at the scattered light on the MC2 frames. The spot is centered on the MC2 approximately.

- MC2 has been aligned so that any resonance is seen in MC_F.

Modification of the MCT optics

- The ND filter before the MCT was removed.

- The Y1-45S mirror before the MCT CCD, which is also used to steer the beam to the MCT QPD path, was replaced to BS1-50-45P.
The reason I used 45P is to obtain higher reflectivity. Because S has higher reflectivity than P in the each layer, I expected to have higher reflectivity for S than 50%.

- The MC REFL path has not been untouched.

Modification of the servo

- The lock was attempted after alignment of the mirrors.  Here how to lock the MC is described below.

1. Run script/MC/mcloopson

2. Open the MC Servo screen in MEDM

3. Change the input gain from 6dB to 22dB.
Change offset from 0.78 to -0.464 (such that the length output has no offset).
Change VCO Gain from 3dB to 21dB

Change MC Length path Gain from 0.3 to 1.6

ELOG V3.1.3-